OpenStudy (anonymous):

ALGEBRA 2 Given f(x) = the quantity of 4x plus 1, divided by 3, solve for f-1(3). A. 2 B. 4 C. 6 D. 8 can somebody explain how to do this please?

OpenStudy (campbell_st):

do you have the inverse function...?

OpenStudy (anonymous):

f(x) = \[\frac{ 4x+1 }{ 3 } solve for f -1 (3) \]

OpenStudy (anonymous):

yes this is a inverse function

OpenStudy (campbell_st):

no, that looks like the function itself.. are you asked to find \[f^{-1}(3)\]

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

I didn't know how to write the inverse on here

OpenStudy (campbell_st):

ok... this is what you need to do swap x and y the original function is \[y = \frac{4x + 1}{3}\] when you make the swap you get \[x = \frac{4y + 1}{3}\] to find the inverse, you need to make y the subject... does that make sense?

OpenStudy (anonymous):

yes this makes sense i've got that far already , im just confused on where to plug in the 3

OpenStudy (campbell_st):

ok... well you can plug it in straight away and say \[3 = \frac{4y + 1}{3}\] and find the value of y

OpenStudy (campbell_st):

or the inverse is \[f^{-1}(x) = \frac{3x -1}{4}\] now substitute x = 3